OpenStudy (ms_sp):

Rachel Levin has a total of $5500 to invest for 1 year. She deposits $4000 into an account earning 6.5% annual simple interest and rest into an account earning 8% annual simple interest. How much interest did Rachel earn?

OpenStudy (ms_sp):

@LannyXX

OpenStudy (opticalmango81):

@Astrophysics

OpenStudy (opticalmango81):

@aly_cat96

OpenStudy (opticalmango81):

@Awolflover1

OpenStudy (opticalmango81):

@TheSmartOne

thomaster (thomaster):

1% is 1 part out of a 100. to get this 1, you can multiply 100 by 0.01, which results in 100. In order to calculate for example 5% of 200, you multiply 200 by 0.05, which is 10. To get 43.5% out of 83,023 you simply multiply 83,023 by 0.435, which is 36,115. So in your case: 4000 * 0.065 = 1500 * 0.08 = _____________________ + total interest earned